OpenStudy (anonymous):

Use the Law of Cosines to solve the triangle to find the length of b, if A= 48degrees, b = 3, and c = 14.

OpenStudy (matt101):

\[a^2=b^2+c^2-2bc \cos A\] \[a^2=3^2+14^2-2(3)(14) \cos 48\] \[a = 12.20\]

OpenStudy (anonymous):

i agree with matt

OpenStudy (matt101):

I'm assuming you wanted to find the length of a, not b.

OpenStudy (matt101):

I agree with DHASHNI :P

OpenStudy (anonymous):

Haha yeah I did. Thanks you guys

OpenStudy (anonymous):

@matt: you foun the length of 'b' and labelled it as 'a' your answer is correct

OpenStudy (matt101):

No dakid91 said b twice in the question, but since angle A and side b were given, I assumed we were going for side a haha

OpenStudy (anonymous):

LOL